I need to sove using Cramer's Rule:
x+3y=21
10x-5y=0

Answer by nerdybill(7384)

x + 3y = 21
10x - 5y = 0
.
coefficient matrix:
1 3
10 -5
.
determinant of coefficient matrix:
(1*-5) - (3*10) = -5 - 30 = -35
.
'x' matrix:
21 3
0 -5
.
determinant of 'x' matrix:
(21*-5) - (0*3) = -105 - 0 = -105
.
'y' matrix:
1 21
10 0
.
determinant of 'y' matrix:
(1*0) - (10*21) = 0 - 210 = -210
.
x = "det of x matrix"/"det of coef matrix"
x = -105/-35 = 3
.
y = "det of y matrix"/"det of coef matrix"
y = -210/-35 = 6
.
Solution: (x,y)=(3,6)